OpenStudy (anonymous):

A random draw is being designed for 210 participants. A single winner is to be chosen, and all the participants must have an equal probability of winning. If the winner is to be drawn using 10 balls numbered 0 through 9, how many balls need to be picked, regardless of order, so that each of the 210 participants can be assigned a unique set of numbers? a-10 b-6 c-5 d-4

OpenStudy (dangerousjesse):

So we're looking for \(\large \frac{x}{10}\)

OpenStudy (dangerousjesse):

Okay, \(10\times 10=100\).. This is one of those questions where they're not asking for the right answer, they want the answer they think is right.

OpenStudy (dangerousjesse):

Oh! Wait, they mean how many combinations of \(x\times 10\)

OpenStudy (dangerousjesse):

100*

OpenStudy (anonymous):

yeah thankssss

OpenStudy (dangerousjesse):

I'd say B, six.

OpenStudy (anonymous):

4 balls are required

OpenStudy (anonymous):

is it correct